### v1

# Symmetry-Resolved Entanglement Entropy from Heat Kernels

Abstract
We develop a systematic framework for computing symmetry-resolved entanglement entropies (SREE) in charged quantum systems based on an improved heat kernel approach. Although the conventional Sommerfeld formula proves effective for neutral systems, it encounters limitations when gauge fields or chemical potentials are introduced due to incomplete residue prescriptions and violations of asymptotic boundary conditions. By reconstructing the analytic structure of the heat kernel using a sign-dependent phase factor, we derive a globally convergent expansion that reconciles discrete residue summations with continuous spectral decompositions. We further apply this framework to Gaussian continuous multi-scale entanglement renormalization ansatz (cMERA) states and show that the entanglement entropy (EE) can be expressed in terms of the cMERA flow functions. In particular, we obtain a symmetry-resolved entanglement entropy flow equation in the presence of a chemical potential. This formulation extends naturally to arbitrary spacetime dimensions and recovers established results for neutral systems in the $\mu\to 0$ limit. We validate our framework through two settings: (1) exact agreement with $(1+1)$-dimensional conformal field theory (CFT) predictions using twist-operator techniques, and (2) consistency with holographic entropy calculations on $S^{1}\times H^{d-1}$ geometries. Our results both unify the treatment of charged and neutral entanglement entropy and extend this treatment to real-space renormalization frameworks, providing a robust tool for probing symmetry-resolved entanglement in conformal field theories, their holographic duals, and cMERA representations.
1 Introduction
Entanglement entropy (EE) has become a cornerstone for understanding emergent phenomena in quantum many-body systems, from critical spin chains to holographic spacetimes. At its core, entanglement entropy quantifies the irreducible correlations between subsystems, a measure that has proven indispensable in characterizing phases of matter, quantum chaos, and the very fabric of spacetime within AdS/CFT correspondence Eisert:2008ur ; Eisert:2006kue ; Horodecki:2009zz ; Levin:2006zz ; Holzhey:1994we ; Calabrese:2009qy ; Calabrese:2004eu ; Calabrese:2016xau ; Calabrese:2009kka ; Cardy:2007mb ; Casini:2005rm ; Casini:2005zv ; Casini:2009sr ; Casini:2011kv ; Ryu:2006bv ; Miyaji:2015yva ; Swingle:2009bg ; Vidal:2007hda ; Vidal:2008zz ; Hung:2011nu ; Balasubramanian:2011wt ; Balasubramanian:2013lsa ; Balasubramanian:2013rqa ; Amico:2007ag ; Nishioka:2018khk ; Latorre:2009zz ; Lewkowycz:2013nqa . However, when one moves beyond the neutral sector toward symmetry-resolved entanglement Belin:2013uta ; Azses:2022nfl ; DiGiulio:2022jjd , fundamental questions naturally arise: How do conserved charges imprint their structure onto entanglement and what mathematical tools can universally capture this interplay between symmetry and quantum correlations?
The interplay between quantum entanglement and global symmetries has emerged as a unifying framework across diverse physical domains, from holographic quantum gravity and topological phases to critical many-body systems. This synergy, driven by novel theoretical constructs Belin:2013uta ; Belin:2014mva ; Pastras:2014oka ; Matsuura:2016qqu ; Laflorencie:2014cem ; Xavier:2018kqb ; Goldstein:2017bua , reveals profound connections between information-theoretic principles and collective quantum phenomena. Importantly, the decomposition of entanglement entropy into distinct symmetry sectors—symmetry-resolved entanglement entropy (SREE)—goes beyond mere theoretical elegance. It encodes the fine-grained quantum correlations constrained by symmetry, thereby offering insights that are inaccessible through total entanglement measures alone. Remarkably, the equipartition, namely, SREE becomes approximately independent of the symmetry sector at leading order, stands as a hallmark property of SREE in gapless systems.
The foundation of SREE lies in quantum information theory, where the Holevo theorem rigorously quantifies the accessible information in symmetry-decomposed mixed states: the maximal extractable information is governed by the entropy difference, which reduces to Shannon entropy in specific limits Xavier:2018kqb . Furthermore, exact computations across diverse theories have established SREE as a universal probe Goldstein:2017bua ; Xavier:2018kqb ; Capizzi:2020jed ; Calabrese:2021wvi ; Bonsignori:2020laa ; Estienne:2020txv ; Ma:2021zgf ; Ares:2022gjb ; Ghasemi:2022jxg ; Murciano:2020vgh ; Horvath:2020vzs ; Horvath:2021fks ; Horvath:2021rjd ; Capizzi:2021kys ; Capizzi:2022jpx ; Capizzi:2022nel ; Foligno:2022ltu ; Fraenkel:2019ykl ; Fraenkel:2021ijv ; Feldman:2019upn ; Murciano:2019wdl ; Murciano:2020lqq ; Murciano:2022lsw ; Calabrese:2020tci ; Turkeshi:2020yxd ; Parez:2020vsp ; Parez:2021pgq ; Ares:2022hdh ; Ares:2022koq ; Piroli:2022ewy ; Scopa:2022gfw . These advances converge on a central insight: the equipartition of SREE emerges as a signature of conformal invariance or integrability, thereby providing a sharp criterion for universality beyond conventional entanglement measures.
In practical simulations, the explicit implementation of symmetry resolution—as manifested in tensor network algorithms such as density matrix renormalization group (DMRG), matrix product state (MPS), and multi-scale entanglement renormalization ansatz (MERA) White1992White ; White:1993zza ; Vidal:2007hda ; Vidal:2008zz ; Verstraete:2008cex —is not merely advantageous but essential. Decomposing the wavefunction into distinct symmetry sectors $\{m\}$ dramatically reduces the resource overhead for high-precision calculations. This efficiency stems from exploiting the block-diagonal structure of symmetric density matrices, enabling studies of complex systems that would otherwise be beyond computational reach.
Recent breakthroughs in quantum simulation have transformed SREE from a purely theoretical construct into a measurable observable. Pioneering experiments with ultra-cold atoms in optical lattices Bloch:2008zzb ; Kaufman:2016mif ; Islam:2015mom ; Lukin:2019tkq ; Azses:2020tdz ; Neven:2021igr ; Vitale:2021lds ; Rath:2022qif have demonstrated the capability to directly probe entanglement. The development of controlled "SWAP" operations between replicated many-body states, as demonstrated with $^{87}\text{Rb}$ atomic chains, provides a concrete pathway for extracting SREE Islam:2015mom ; Lukin:2019tkq ; Azses:2020tdz ; Vitale:2021lds ; Neven:2021igr ; Rath:2022qif . This methodology, which bridges quantum information protocols with AMO platforms, opens unprecedented avenues for the experimental validation of symmetry-resolved quantum correlations.
Traditional approaches to SREE are primarily based on twist operator correlators in conformal field theory (CFT) Murciano:2020vgh and minimal surface prescriptions in holography Belin:2013uta ; Huang:2025lsy . The heat kernel method, owing to its geometric intuition and adaptability to replica manifolds Solodukhin:2011gn ; Dowker:1977zj ; Fursaev:1994in ; Mann:1997hm ; Frolov:1998ea , offers an alternative approach.
For a quantum field $\Phi$ governed by a Gaussian Lagrangian $\mathcal{L}=\Phi^{\dagger}D\Phi$, the effective action $W[\Phi]$ can be expressed in terms of the heat kernel $K(s,x,x^{\prime})$, which solves the heat kernel equation associated with the operator $D$:
$W[\Phi]=-\frac{1}{2}\int_{0}^{\infty}\frac{ds}{s}\operatorname{Tr}K(s).$ (1)
In standard replica calculations Calabrese:2004eu ; Calabrese:2009qy ; Calabrese:2016xau , boundary conditions on $n$-sheeted Riemann surfaces $\mathcal{R}_{n}$ are systematically handled using the Sommerfeld formula Fursaev:1994in ; Dowker:1977zj , which adjusts the periodicity of field configurations to account for topological defects. This approach has proven effective for neutral systems, in which the analytic structure of $\operatorname{Tr}K(s)$ remains well controlled.
However, the introduction of background $U(1)$ gauge fields $A_{\mu}(x)$ in charged quantum systems reveals significant limitations in the conventional Sommerfeld framework. Specifically: (i) Incomplete residue analysis: Conventional Sommerfeld formulas account only for singularities arising from insertion functions while neglecting additional singularities induced by the non-analytic nature of heat kernels, thus leading to systematic discrepancies between residue summations and discrete Sommerfeld expansions; (ii) Breakdown of asymptotic boundary conditions: Contour integral treatments lack rigorous justification for neglecting contributions from infinite boundaries, an approximation that fails when the chemical potential $\mu\neq 0$.
To address these challenges, this work develops a systematic approach that reconciles the analytic structure of heat kernels with the constraints imposed by a finite chemical potential. Our approach reconstructs the heat kernel representation of the effective action through a careful reformulation of the analytic structure of Sommerfeld formulas. The main innovations are: (i) replacement of imposed periodic boundary conditions by asymptotic decay constraints; (ii) introducing a kernel function endowed with a sign-dependent phase factor, which systematically handles charged configurations and eliminates contributions from infinite contours. This refinement enables the first globally convergent expansion for charged heat kernels on $n$-sheeted manifolds, thus bridging the gap between discrete Sommerfeld-type expansions and continuous spectral decompositions.
We obtain explicit expressions for the $d$-dimensional charged Rényi entropy of scalar fields. These results not only generalize existing formulas for neutral systems, but also recover traditional Sommerfeld predictions in the $\mu\to 0$ limit, thereby validating the consistency of our approach.
Furthermore, we verify the universality of our framework through two independent benchmarks. First, for (1+1)(1+1)D CFT models, the SREE calculated using our heat kernel approach agrees exactly with the results derived from twist operator techniques in Murciano:2020vgh . Second, computations of entanglement entropy for conformal field theories on $S^{1}\times H^{d-1}$ backgrounds match the analytic predictions from Belin:2013uta ; Huang:2025lsy . Together, these validations confirm that the improved heat-kernel method remains robust across both low-dimensional conformal systems and higher-dimensional geometric settings.
Finally, building on the known cMERA representation of neutral entanglement entropy Fernandez-Melgarejo:2021ymz , we extend the surface contribution of the replicated heat kernel to Gaussian cMERA states in the presence of a chemical potential. In particular, we derive a modified cMERA entanglement flow equation in the charged case and demonstrate its smooth reduces to the neutral result as $\mu\to 0$.
The paper is organized as follows. Section 2 reviews the heat kernel representation of effective actions and conventional Sommerfeld expansion. Section 3 presents the improved Sommerfeld scheme featuring the sign function kernel, which is validated in $d$-dimensional neutral free field theories. In Section 4 this scheme is applied to compute the symmetry-resolved entanglement entropy for free CFTs in flat $d$-dimensional spacetime as well as on $S^{1}\times H^{d-1}$ backgrounds. Section 5 establishes the correspondence between the surface heat kernel and Gaussian cMERA correlation functions, and derives a modified entanglement flow equation in the presence of a chemical potential. Finally, Section 6 summarizes the broader implications of the framework and outlines potential extensions to MERA and holographic duality.
2 Entanglement and Heat Kernels
2.1 Entanglement Entropy and Symmetry Resolution
[…]
$S$ $\equiv-\operatorname{Tr}(\rho_{A}\ln\rho_{A}),$ (2)
$S_{n}$ $\equiv\frac{1}{1-n}\ln\operatorname{Tr}\rho_{A}^{n},$ (3)
with the relation $S=\lim_{n\to 1}S_{n}$ obtained through analytic continuation from integer values of $n$ to real values. These entropies can be computed using the replica trick, which involves $n$-sheeted Riemann surfaces $\mathcal{R}_{n}$ and topological defects, by taking derivatives of the effective action with respect to the replica parameter Calabrese:2004eu . Specifically, on an $n$-sheeted Riemann surface, we have:
$\operatorname{Tr}\rho^{n}_{A}=\frac{Z_{n}}{Z_{1}^{n}},$ (4)
where $Z_{n}\equiv Z[\mathcal{R}_{n}]$ denotes the partition function incorporating topological defects. Notably, the Rényi entropy can be reformulated as
$S_{n}=\frac{1}{1-n}\left(\ln Z_{n}-n\ln Z_{1}\right),$ (5)
as shown in Calabrese:2004eu . By analytic continuation, we express $S$ as
$S=-\left.\frac{\partial}{\partial\alpha}\ln\frac{Z_{\alpha}}{Z_{1}^{\alpha}}\right|_{\alpha=1}=\left.\left(\alpha\frac{\partial}{\partial\alpha}-1\right)W_{\alpha}\right|_{\alpha=1},$ (6)
where the effective action $W_{\alpha}$ is defined via the replicated partition function:
$W_{\alpha}\equiv-\ln Z_{\alpha}+\alpha\ln Z_{1}.$ (7)
Here, the term $\alpha\ln Z_{1}$ ensures the cancellation of boundary contributions from individual replicas. For normalized vacuum states ($Z_{1}=1$), this simplifies to
$S=-\left.\partial_{\alpha}W_{\alpha}\right|_{\alpha=1}.$ (8)
This framework achieves remarkable success in conformal field theory. For $(1+1)$-dimensional CFTs on an infinite line, the Calabrese-Cardy method provides analytic solutions for the Rényi entropy of a single interval of length $l$ through twist operator correlation functions
$S_{n}=\frac{c}{6}\left(1+\frac{1}{n}\right)\ln\left(\frac{l}{\epsilon}\right),\qquad S=\frac{c}{3}\ln\left(\frac{l}{\epsilon}\right),$ (9)
where $c$ is the central charge, $l$ is the interval length, and $\epsilon$ is the ultraviolet cutoff. For finite systems of total size $R$ with periodic boundary conditions (e.g., a circle), the entropy for half the system ($l=R/2$) becomes:
$S_{n}=\frac{c}{12}\left(1+\frac{1}{n}\right)\ln\left(\frac{R}{\pi\epsilon}\right),\qquad S=\frac{c}{6}\ln\left(\frac{R}{\pi\epsilon}\right),$ (10)
reflecting the topological constraints imposed by the compactified spatial dimension Calabrese:2004eu ; Calabrese:2009qy ; Calabrese:2016xau .
A groundbreaking advancement emerges within the AdS/CFT correspondence tHooft:1993dmi ; Susskind:1994vu ; Maldacena1999 ; Witten:1998qj ; Gubser:1998bc : The Ryu-Takayanagi formula maps holographic entanglement entropy to minimal surface areas in AdS spacetime Ryu:2006bv ,
$S=\frac{\text{Area}(\gamma)}{4G_{N}},$ (11)
revealing an intrinsic connection between quantum entanglement and spacetime geometry. In particular, the continuous Multi-scale Entanglement Renormalization Ansatz (cMERA) Haegeman:2011uy —a field-theoretic realization of tensor networks—exhibits a profound correspondence between its renormalization group flow and AdS geometry Swingle:2009bg ; Nozaki:2012zj . This correspondence provides powerful new tools for probing the microscopic mechanisms of holography via refined entanglement entropy frameworks.
Recent extensions of this framework to SREE have yielded substantial progress. Consider a system with a conserved $U(1)$ charge generated by an operator $Q$. The Hilbert space decomposes naturally into distinct charge sectors labeled by the eigenvalues $q$. The reduced density matrix $\rho_{A}$ can be block-diagonalized as
$\rho_{A}=\oplus_{q}p_{A}(q)\rho_{A}(q),\quad p_{A}(q)=\operatorname{Tr}(\Pi_{q}\rho_{A}),$ (12)
where $\Pi_{q}$ is the projection operator onto the charge-$q$ subspace, and $\rho_{A}(q)$ describes the normalized density matrix within this sector. The symmetry-resolved Rényi entropy for sector $q$ is then defined as
$S_{n}(q)\equiv\frac{1}{1-n}\ln\left[\frac{Z_{n}(q)}{Z_{1}^{n}(q)}\right],\quad Z_{n}(q)=\operatorname{Tr}\left(\Pi_{q}\rho_{A}^{n}\right).$ (13)
where $Z_{n}(q)$ are the symmetry-resolved moments. They are related to the charged moments $Z_{n}(\mu)$ through a Fourier transform:
$Z_{n}(q)=\int_{-\pi}^{\pi}\frac{d\mu}{2\pi}\,e^{-iq\mu}Z_{n}(\mu),$ (14)
so that $q$ and $\mu$ play conjugate roles as the discrete charge and its associated chemical potential.
To compute $Z_{n}(q)$, we introduce a fixed background $U(1)$ gauge field that couples to the conserved current Belin:2013uta . This background field is flat ($dA=0$) everywhere except at the entangling surface $\Sigma$, where it carries a nontrivial Wilson line $\oint_{C}A=-in\mu$.
The charged Rényi moments $Z_{n}(\mu)$ are naturally expressed as a partition function in the replicated manifold $\mathcal{R}_{n}$, with the gauge field $A_{\mu}(x)$ introducing twisted boundary conditions across replicas. This yields:
$Z_{n}(\mu)=\operatorname{Tr}\left[\rho_{A}e^{\mu Q_{A}}\right]^{n},$ (15)
and the corresponding charged Rényi entropy reads
$S_{n}(\mu)=\frac{1}{n-1}\left(F_{n}(\mu)-nF_{1}(\mu)\right),$ (16)
where $F_{n}(\mu)=-\ln Z_{n}(\mu)$ denotes the corresponding free energy.
2.2 Heat Kernel Representation for Effective Actions
The heat kernel method provides a systematic framework for computing effective actions and entanglement entropy in quantum field theory. Its particular effectiveness in curved spacetime backgrounds makes it a critical tool for evaluating entanglement entropy in nontrivial geometries, particularly the conical geometries that emerge in conformal field theory applications. The foundational analyses by Fursaev and Frolov Fursaev:1994in ; Frolov:1996aj comprehensively address its implementation on manifolds containing conical singularities.
For a bosonic field $\Phi$, the effective action $W$ under Gaussian approximation is fundamentally expressed as:
$e^{-W}=(\det D)^{-1/2},\quad W=\frac{1}{2}\operatorname{Tr}\ln D,$ (17)
where $D$ is a positive-definite elliptic operator acting on the field space. In this work, unless otherwise specified, $D$ will be taken to be the Laplace operator. Assume that $D$ admits a complete orthonormal eigenbasis:
$D|d_{n}\rangle=d_{n}|d_{n}\rangle,\quad\langle d_{m}|d_{n}\rangle=\delta_{mn}.$ (18)
The determinant and trace-logarithm then follow directly:
$\det D=\prod_{n}d_{n},\quad\operatorname{Tr}\ln D=\sum_{n}\ln d_{n}.$ (19)
Using the integral identity:
$\ln d_{n}=-\int_{0}^{\infty}\frac{ds}{s}e^{-sd_{n}},$ (20)
the effective action transforms into:
$W=-\frac{1}{2}\int_{0}^{\infty}\frac{ds}{s}\sum_{n}e^{-sd_{n}}.$ (21)
The corresponding heat kernel is defined as
$K(s;x,x^{\prime})\equiv\langle x|e^{-sD}|x^{\prime}\rangle=\sum_{n}e^{-sd_{n}}\braket{x|d_{n}}\braket{d_{n}|x^{\prime}},$ (22)
satisfies the heat equation with initial condition:
$(\partial_{s}+D)K(s;x,x^{\prime})=0,\quad K(0;x,x^{\prime})=\delta^{(d)}(x-x^{\prime}).$ (23)
The heat kernel trace evaluates to:
$\operatorname{Tr}K(s)=\int dx\,\ K(s;x,x)=\sum_{n}e^{-sd_{n}},$ (24)
yielding the final expression for $W$:
$W=-\frac{1}{2}\int_{0}^{\infty}\frac{ds}{s}\operatorname{Tr}K(s).$ (25)
2.3 Sommerfeld Formula
In flat space, the Sommerfeld formula provides an efficient approach for computing heat kernel contributions from singular surfaces $\Sigma$. This construction admits generalization to arbitrary curved spaces $E_{\alpha}$ that possess at least one local Abelian isometry with a fixed point Solodukhin:2011gn .
The strength of the Sommerfeld method lies in its ability to systematically handle modified periodicity requirements. For Rényi entropy calculations through replica trick approaches, the heat kernel must satisfy $2\pi\alpha$-periodic boundary conditions induced by conical geometries. The Sommerfeld formalism reconstructs the heat kernel satisfying modified periodicity by representing it as a discrete sum or, equivalently, as a contour integral.
Furthermore, this method naturally extends to systems with phase-dependent boundary conditions that arise from chemical potentials or gauge connections. Through its representation as either discrete sums or contour integrals, the Sommerfeld-modified heat kernel automatically incorporates these generalized boundary conditions. These properties establish the combined heat kernel-Sommerfeld approach as particularly effective for entropy calculations in non-trivial geometric backgrounds.
Consider a heat kernel $K(s;\phi,\phi^{\prime})$ that satisfies standard $2\pi$-periodic boundary conditions, where $\phi$ denotes an angular coordinate:
$K(s;\phi-\phi^{\prime})=K(s;\phi-\phi^{\prime}+2\pi).$ (26)
To adapt this heat kernel for Rényi entropy computations that require $2\pi\alpha$-periodicity, we implement the Sommerfeld modification through a correction term $\Delta_{\alpha}$:
$K_{\alpha}(s;\phi,\phi^{\prime})=K(s;\phi,\phi^{\prime})+\Delta_{\alpha}(s;\phi,\phi^{\prime}).$ (27)
The modified kernel must then satisfy:
$K_{\alpha}(s;\phi-\phi^{\prime})=K_{\alpha}(s;\phi-\phi^{\prime}+2\pi\alpha).$ (28)
Following the standard construction in the literature Dowker:1977zj ; Fursaev:1994in ; Solodukhin:2011gn , this modified kernel $K_{\alpha}(s;\phi,\phi^{\prime})$ is obtained from the periodic one through the Sommerfeld formula:
$K_{\alpha}(s;\phi,\phi^{\prime})=K(s;\phi,\phi^{\prime})+\frac{i}{4\pi\alpha}\int_{\Gamma_{0}}\cot\left(\frac{w}{2\alpha}\right)K(s;\phi-\phi^{\prime}+w)dw,$ (29)
where the contour $\Gamma_{0}$ consists of two vertical segments: $(-\pi+i\infty)\to(-\pi-i\infty)$ and $(\pi-i\infty)\to(\pi+i\infty)$, as illustrated in Fig. 1. The real parts $\operatorname{Re}w=\pm\pi$ are chosen so that the contour encloses all poles of the integrand except the one at $w=0$. These poles are located at $w=2\pi\alpha m(m\in\mathbb{Z})$, where $\alpha$ is the replica parameter that will be continued to values close to $1$; thus, the choice $\pm\pi$ naturally captures all nonzero poles within a $2\pi$-wide fundamental strip. The normalization factor originates from the residue calculation:
$2\pi i\;\mathrm{Res}\left(\cot\left(\frac{w}{2\alpha}\right),w=0\right)=4\pi i\alpha.$ (30)
Figure 1: Integration contours $\Gamma_{0}$, $\Gamma_{1}$, and $\Gamma_{2}$ in the complex $w$-plane. All contours are topologically equivalent, enclosing the same set of poles except for the one at the origin.
3 Sommerfeld Formula for Charged Systems
3.1 Modified Phase Factors and Convergence
Within the framework of quantum field theory, computing entanglement entropy via the replica trick involves the $n$-th power of the reduced density matrix, expressed through the Euclidean path integral $\operatorname{Tr}\rho^{n}=Z[\mathcal{R}_{n}]$ on the $n$-fold branched cover $\mathcal{R}_{n}$ of the cut spacetime manifold.
This corresponds geometrically to a conical geometry with deficit angle $2\pi(1-\alpha)$ localized at the entangling surface $\Sigma$, where the analytic continuation from $n$ to non-integer values $\alpha$ is implicit. The entanglement entropy emerges through the variational identity (6):
$S=\left.\left(\alpha\frac{\partial}{\partial\alpha}-1\right)W(\alpha)\right|_{\alpha=1}.$ (31)
The aim of this section is to generalize the Sommerfeld construction of the heat kernel to charged configurations. While the conventional Sommerfeld formula Solodukhin:2011gn suffices for neutral cases, we present a refined version that systematically accommodates charged fields while preserving geometric intuition. The heat kernel is a solution of the heat kernel equation, and the essence of the Sommerfeld method is to construct new solutions by linearly superposing basis kernels so as to impose the desired boundary conditions. For the conical geometry with opening angle $2\pi\alpha$, this can be realized by series summation:
$K_{\alpha}(s;\phi,\phi^{\prime})=\sum_{m\in\mathbb{Z}}K\big(s;\phi-\phi^{\prime}+2\pi m\alpha\big),$ (32)
which enforces the periodicity condition Eq. (28) .
To transform the discrete representation Eq. (32) into a continuous one, we employ contour integration in the complex $w$-plane. The idea is to replace the discrete index $m$ by a set of poles of an auxiliary insertion function, so that the discrete sum is recovered through residue calculus. This yields the standard (continuous) Sommerfeld representation Eq. (29) .
However, a closer inspection reveals two subtleties in Eq. (29). First, the residue calculus in Eq. (29) must simultaneously account for both the poles of $\cot(w/2\alpha)$ at $w=2\pi m\alpha$ and the singularities of the heat kernel $K(s;\phi,\phi^{\prime})$. If $K(s;\phi,\phi^{\prime})$ were assumed to be periodic in $\phi-\phi^{\prime}$, its singularities would be periodically replicated, and the contour deformation linking the discrete and continuous formulations would become ill-defined. This mismatch leads to discrepancies between discrete and continuum contributions.
Second, in the charged case ($\mu\neq 0$), we should further consider the potential divergence contributed by the complex infinity in Eq. (29). We consider the situation where the chemical potential is purely imaginary, $\mu=i\mu_{E}$ with real $\mu_{E}$. Consequently, the heat kernel acquires a phase factor $e^{-iw\mu_{E}/(2\pi\alpha)}$, so that the integrand behaves as
$I(w)\sim e^{-i\frac{w}{2\pi\alpha}\mu_{E}}\,\cot\frac{w}{2\alpha}\,K(s;\phi-\phi^{\prime}+w).$ (33)
Writing $w=u+iv$, the phase factor decomposes into
$e^{-i\frac{w}{2\pi\alpha}\mu_{E}}=e^{-i\frac{u}{2\pi\alpha}\mu_{E}}\,e^{\frac{v}{2\pi\alpha}\mu_{E}}.$ (34)
Hence, for $\mu_{E}>0$, the integrand grows exponentially as $v\to+\infty$ and decays as $v\to$-$\infty$; the opposite holds for $\mu_{E}<0$. Consequently, the contour integral diverges along one side of the imaginary axis unless a regulator is introduced.
Our resolution involves two key modifications of the formalism. First, since a consistent residue analysis requires that the heat kernel $K(s;\phi-\phi’+w)$ be analytic on the real $w$-axis except at the origin, we impose this condition as a basic assumption. To realize it, we no longer require the heat kernel $K(s;\phi,\phi^{\prime})$ to satisfy periodic boundary conditions Eq. (26). Equivalently, we assume infinite periodicity to avoid introducing unnecessary singularities. Second, we replace the insertion function $\cot(w/2\alpha)$ with the regulator function, designed to regularize the large-imaginary-$w$ behavior:
$\Xi(w)\equiv\frac{1}{1-e^{-i\frac{w}{\alpha}\mathrm{sgn}(\mu_{E})}},$ (35)
where $\mathrm{sgn}[x]$ denotes the sign function and $\mu_{E}$ denotes the Euclidean chemical potential along the compact imaginary-time direction. In the neutral case, this term vanishes, whereas in the charged case it maintains poles at $w=2\pi\alpha m$ but provides exponential damping at large $|\mathrm{Im}\,w|$.
The geometric interpretation in Fig. 1 clarifies the role of the regulator. The contour $\Gamma_{1}$ is a closed path consisting of large arcs in the upper and lower half-planes, bending inward along the real axis to enclose all singularities on the real axis except the origin. The regulator $\Xi(w)$ ensures that, as $|w|\to\infty$ along either semicircle, the integrand vanishes, making the contribution from $\Gamma_{1}$ identically zero. This directly illustrates why the sign function in $\Xi(w)$ guarantees convergence at infinity. Meanwhile, $\Gamma_{2}$ is a small clockwise circle around the origin; contributions from all other singularities are captured by taking the residue at $w=0$ with opposite sign.
The improved Sommerfeld formula takes the form
$K_{\alpha\mu}(s;\phi,\phi^{\prime})=K(s;\phi,\phi^{\prime})+\frac{\mathrm{sgn}(\mu_{E})}{2\pi\alpha}\int_{\Gamma_{2}}e^{-i\frac{w}{2\pi\alpha}\mu_{E}}\frac{1}{1-e^{-i\frac{w}{\alpha}\mathrm{sgn}(\mu_{E})}}K(s;\phi-\phi^{\prime}+w)dw,$ (36)
where the normalization is fixed by the residue
$2\pi i\;\mathrm{Res}\left(\frac{1}{1-e^{-i\frac{w}{\alpha}\mathrm{sgn}(\mu_{E})}},w=0\right)=2\pi\alpha\;\mathrm{sgn}(\mu_{E}).$ (37)
The regulator $\Xi(w)$ preserves the periodic poles, cancels the exponential divergence introduced by the charged phase, and ensures convergence of the contour integral.
3.2 Consistency in Polar Coordinates
We now verify the applicability of the improved Sommerfeld formula to $d$-dimensional free bosonic fields. To this end, we first solve the heat kernel and its trace in polar coordinates before applying the replica trick to evaluate the entanglement entropy.
3.2.1 Heat Kernel on Multi-sheeted Riemann Surfaces
Consider a $d$-dimensional spacetime containing a (d2)(d-2)-dimensional entangling surface. The solution to the heat equation Eq. (23) obtained by Fourier transformation reads
$K(s;x,x^{\prime})=\frac{1}{(2\pi)^{d}}\int d^{d}p\;e^{ip_{\mu}(x^{\mu}-x^{\prime\mu})}e^{-sp^{2}}.$ (38)
To evaluate the trace of the heat kernel $K(s;x,x^{\prime})$, we first set the coordinates equal on the (d2)(d-2)-dimensional hyperplane: $z_{i}=zi$ for $i=1,\dots,d-2$. In the remaining two dimensions, we adopt polar coordinates $(r,\phi)$ for both points and take $r=r^{\prime}$, while defining the angular difference as $w=\phi-\phi^{\prime}$. With this choice, the distance becomes $|x-x^{\prime}|=2r\sin(w/2)$. Let $p=|p_{\mu}|$ and $\theta$ be the angle between $p_{\mu}$ and $(x-x^{\prime})_{\mu}$, then
$p_{\mu}(x^{\mu}-x^{\prime\mu})=2pr\sin\!\left(w/2\right)\cos\theta.$ (39)
Since in our present formulation the heat kernel is regarded as non-periodic in the angular variable, unlike the standard periodic construction on a cone, it is no longer necessary to preserve the exact trigonometric form of $\sin(w/2)$. Accordingly, we may expand it around $w=0$ :
$\sin(w/2)=\frac{w}{2}-\frac{w^{3}}{48}+O(w^{4}),$ (40)
and retain the two leading orders. This truncation also gets rid of the zeros $\omega=2\pi n,\,\forall n\in\mathbb{Z}$ and $n\neq 0$, which is consistent with our assumption that the heat kernel is analytic on the real axis except at the origin. Substituting this expansion gives
$p_{\mu}(x^{\mu}-x^{\prime\mu})\simeq 2pr\left(\frac{w}{2}-\frac{w^{3}}{48}\right)\cos\theta,$ (41)
which will be sufficient for the subsequent small-$w$ analysis.
The integration measure for the $d$-dimensional momentum $p_{\mu}$ in spherical coordinates reads:
$\int d^{d}p=\Omega_{d-2}\int p^{d-1}dp\int\sin^{d-2}\theta d\theta,$ (42)
where $\Omega_{d-2}=2\pi^{(d-1)/2}/\Gamma[(d-1)/2]$. Substituting into Eq. (38) yields:
$K(s;w,r)=\frac{\Omega_{d-2}}{(2\pi)^{d}}\int_{0}^{\infty}\!p^{d-1}\,dp\int_{0}^{\pi}\!\sin^{d-2}\!\theta\,e^{i\,2pr\left(\frac{w}{2}-\frac{w^{3}}{48}\right)\cos\theta}e^{-sp^{2}}d\theta.$ (43)
The angular integral can be simplified using the Bessel function identity:
$\int_{0}^{\pi}\sin^{\nu}\theta e^{ix\cos\theta}d\theta=\sqrt{\pi}\,{\Gamma\left(\frac{\nu+1}{2}\right)}\left(\frac{2}{x}\right)^{\nu/2}J_{\nu/2}(x),$ (44)
where $J_{\nu/2}(x)$ is the Bessel function of the first kind, we obtain
$K(s;w,r)=\frac{2}{(4\pi)^{d/2}}\frac{1}{\bigl[r\,(\frac{w}{2}-\frac{w^{3}}{48})\bigr]^{\frac{d-2}{2}}}\int_{0}^{\infty}dp\,p^{d/2}\,J_{\frac{d-2}{2}}\Bigl(2pr\,(\frac{w}{2}-\frac{w^{3}}{48})\Bigr)\,e^{-sp^{2}}.$ (45)
The full trace over configuration space requires integration over $r$, $\phi$, and transverse coordinates:
$\operatorname{Tr}K(s;w)$ $=\int d^{d-2}z\int_{0}^{2\pi}d\phi\int_{0}^{\infty}rdr\,K(s;w,r)$ (46)
$=\frac{s}{(4\pi s)^{d/2}}\cdot\pi\alpha A(\Sigma)\cdot\left(\frac{4}{w^{2}}+\frac{1}{3}\right),$ (47)
where $A(\Sigma)=\int d^{d-2}z$ is the entangling surface area. This result agrees with the known expressions in Solodukhin:2011gn .
Application of the improved Sommerfeld formula Eq. (36) modifies the trace through
$\operatorname{Tr}\Delta_{\alpha}=\operatorname{Tr}\left[\frac{1}{2\pi\alpha}\int_{\Gamma_{2}}\frac{1}{1-e^{-iw/\alpha}}K(s;\phi-\phi^{\prime}+w)dw\right].$ (48)
Substituting our trace expression Eq. (47) produces
$\operatorname{Tr}\Delta_{\alpha}$ $=\frac{s}{(4\pi s)^{d/2}}\frac{A(\Sigma)}{2}\int_{\Gamma_{2}}\frac{1}{1-e^{-i\frac{w}{\alpha}}}\left(\frac{4}{w^{2}}+\frac{1}{3}\right)dw$ (49)
$=\frac{1}{(4\pi s)^{(d-2)/2}}A(\Sigma)\cdot\frac{\alpha C_{2}(\alpha)}{2},$ (50)
where
$C_{2}(\alpha)$ $=\frac{1}{4\pi\alpha}\int_{\Gamma_{2}}\frac{1}{1-e^{-i\frac{w}{\alpha}\,}}\left(\frac{4}{w^{2}}+\frac{1}{3}\right)dw.$ (51)
As the contour $\Gamma_{2}$ encloses all poles except the origin, only the residue at $w=0$ contributes:
$\mathrm{Res}\left(\frac{1}{1-e^{-i\frac{w}{\alpha}\,}}\left(\frac{4}{w^{2}}+\frac{1}{3}\right),w=0\right)$ $=\frac{i\left(\alpha^{2}-1\right)}{3\alpha}.$ (52)
yielding
$C_{2}(\alpha)$ $=\frac{1}{6}\left(\frac{1}{\alpha^{2}}-1\right).$ (53)
3.2.2 Entanglement Entropy from Heat Kernel
The effective action can be evaluated as
$W[\alpha]$ $=-\frac{1}{2}\int_{\epsilon^{2}}^{\infty}\frac{ds}{s}\operatorname{Tr}K_{\alpha}(s)$ (54)
$=-\frac{1}{2}\int_{\epsilon^{2}}^{\infty}\frac{ds}{s}\left[\frac{1}{(4\pi s)^{d/2}}V\cdot\alpha+\frac{1}{(4\pi s)^{(d-2)/2}}A(\Sigma)\cdot\frac{\alpha C_{2}(\alpha)}{2}\right]$ (55)
Applying the entropy formula Eq. (6), we obtain
$S=\frac{A(\Sigma)}{6(d-2)(4\pi)^{(d-2)/2}\epsilon^{d-2}}.$ (56)
In $d=2$ , dimensional regularization leads to logarithmic scaling:
$S_{d=2}=\frac{\ln\epsilon}{6}A(\Sigma).$ (57)
This consistency in neutral systems provides the foundation for extending the formalism to charged fields, where the sign-dependent phase factor becomes essential as developed in Section 3.1.
4 Heat Kernel Method for SREE
We develop a generalized heat kernel formalism for charged systems by incorporating a finite chemical potential $\mu_{E}$. This extension provides a unified computational framework for both the free energy and entanglement entropy in the presence of conserved charges.
4.1 Heat Kernel for Charged Rényi Entropies
Building upon the improved Sommerfeld formalism with sign-dependent phase factors developed in Section 3, we now extend the heat kernel framework to compute charged Rényi entropies. The key modification lies in imposing phase-twisted boundary conditions that encode the chemical potential dependence. In this framework, the heat kernel satisfies the phase boundary conditions as specified in Eq. (A.2) of Belin:2013uta :
$K_{\alpha\mu}(s;\phi-\phi^{\prime}+2\pi m\alpha)=e^{im\mu_{E}}K_{\alpha\mu}(s;\phi-\phi^{\prime}),$ (58)
leading to the discrete spectral representation:
$K_{\alpha\mu}(s;\phi,\phi^{\prime})=\sum_{m=-\infty}^{\infty}e^{-im\mu_{E}}K(s;\phi-\phi^{\prime}+2\pi m\alpha).$ (59)
The corresponding continuous version is:
$K_{\alpha\mu}(s;\phi,\phi^{\prime})=K(s;\phi,\phi^{\prime})+\frac{i}{4\pi\alpha}\int_{\Gamma_{2}}e^{-i\frac{w}{2\pi\alpha}\mu_{E}}\frac{1}{1-e^{-i\frac{w}{\alpha}\,\mathrm{sgn}(\mu_{E})}}K(s;\phi-\phi^{\prime}+w)dw.$ (60)
Under the phase condition, the trace of the Sommerfeld correction term can be computed similarly to equation (48) as:
$\operatorname{Tr}\Delta_{\alpha\mu}$ $=\operatorname{Tr}\left[\frac{\mathrm{sgn}(\mu_{E})}{2\pi\alpha}\int_{\Gamma_{2}}e^{-i\frac{w}{2\pi\alpha}\mu_{E}}\frac{1}{1-e^{-i\frac{w}{\alpha}\,\mathrm{sgn}(\mu_{E})}}K(s;\phi-\phi^{\prime}+w)dw\right].$ (61)
4.2 SREE for Free Scalar Fields
We now apply this framework to free scalar field theories, utilizing the heat kernel trace expression derived in Eq. (47). The central quantity of interest is the coefficient function:
$C_{2}(\alpha,\mu_{E})$ $=\frac{\mathrm{sgn}(\mu_{E})}{4\alpha\pi}\int_{\Gamma_{2}}e^{-i\frac{w}{2\pi\alpha}\mu_{E}}\frac{1}{1-e^{-i\frac{w}{\alpha}\,\mathrm{sgn}(\mu_{E})}}\left(\frac{4}{w^{2}}+\frac{1}{3}\right)dw.$ (62)
The residue calculation at $w=0$ yields the essential singularity structure:
$\mathrm{Res}\left(e^{-i\frac{w}{2\pi\alpha}\mu_{E}}\frac{1}{1-e^{-i\frac{w}{\alpha}\,\mathrm{sgn}(\mu_{E})}}\left(\frac{4}{w^{2}}+\frac{1}{3}\right),w=0\right)$
$=-\frac{i\mu_{E}}{\pi\alpha}+\frac{i\mu_{E}^{2}}{2\pi^{2}\alpha\text{sgn}(\mu_{E})}+\frac{i\text{sgn}(\mu_{E})}{3\alpha}-\frac{i\alpha}{3\text{sgn}(\mu_{E})}$ (63)
Consequently, the coefficient function takes the form:
$C_{2}(\alpha,\mu_{E})=\frac{1}{6}\left(\frac{1}{\alpha^{2}}-1\right)+\frac{\mu_{E}^{2}}{4\pi^{2}\alpha^{2}}-\frac{|\mu_{E}|}{2\pi\alpha^{2}}.$ (64)
For charged systems, the charged Rényi entropy is expressed in terms of the free energy $F_{n}(\mu_{E})$ as:
$S_{n}(\mu_{E})=\frac{F_{n}(\mu_{E})-nF_{1}(\mu_{E})}{n-1},$ (65)
where the free energy can be computed via the heat kernel trace as
$F_{n}(\mu_{E})$ $=-\frac{1}{2}\int_{\epsilon^{2}}^{\infty}\frac{ds}{s}\operatorname{Tr}K_{n\mu}(s)$ (66)
$=-\frac{n\left[(d-2)V+2\pi d\,\epsilon^{2}\,A(\Sigma)\,C_{2}(n,\mu_{E})\right]}{d(d-2)(4\pi)^{d/2}\,\epsilon^{d}}.$ (67)
Thus, the Rényi entropy is
$S_{n}(\mu_{E})=-\frac{A(\Sigma)}{2(d-2)(4\pi)^{(d-2)/2}\epsilon^{d-2}}\frac{n\left[C_{2}(n,\mu_{E})-nC_{2}(1,\mu_{E})\right]}{n-1}.$ (68)
Taking the analytic continuation $n\to 1$, we obtain the $\mu_{E}$-dependent entanglement entropy:
$S(\mu_{E})=\frac{A(\Sigma)}{(d-2)(4\pi)^{(d-2)/2}\epsilon^{d-2}}\left(\frac{1}{6}-\frac{|\mu_{E}|}{2\pi}+\frac{\mu_{E}^{2}}{4\pi^{2}}\right).$ (69)
For the special case of $d=2$ dimensions, this expression reduces to logarithmic scaling:
$S(\mu_{E})=-\frac{\ln(\epsilon)}{2}A(\Sigma)\left(\frac{|\mu_{E}|}{\pi}-\frac{\mu_{E}^{2}}{2\pi^{2}}-\frac{1}{3}\right),$ (70)
matching known CFT results Azses:2022nfl .
4.3 SREE for $S^{1}\times H^{d-1}$ Background
Following the framework developed in Belin:2013uta , we now examine the heat kernel structure and free energy on $S^{1}\times H^{d-1}$ geometries. This analysis extends our formalism to curved backgrounds relevant for holographic entanglement entropy calculations.
Heat Kernel Factorization on $S^{1}\times H^{d-1}$
The product geometry of $S^{1}\times H^{d-1}$ enables factorization of the heat kernel:
$K_{S^{1}\times H^{d-1}}(s;x,y)=K_{S^{1}}(s;x_{1},y_{1})K_{H^{d-1}}(s;x_{2},\dots,x_{d},y_{2},\dots,y_{d}),$ (71)
where $K_{S^{1}}$ and $K_{H^{d-1}}$ are the heat kernels on $S^{1}$ and $H^{d-1}$, respectively.
The corresponding free energy on $S^{1}\times H^{d-1}$ can be expressed as an integral of the heat kernel:
$F=-(-1)^{f}\frac{1}{2}\int\frac{ds}{s}\operatorname{Tr}K_{S^{1}\times H^{d-1}}(s)e^{-m_{c}(1-f)s},$ (72)
where $f$ encodes the particle statistics ($f=1$ for Dirac fermions, $f=0$ for scalars) and $m_{c}=-\frac{(d-2)^{2}}{4R^{2}}$ represents the conformal coupling mass with $R=1$.
The charged system imposes phase-twisted boundary conditions on the heat kernel:
$K_{S^{1}\times H^{d-1}}(s;x_{1}+2\pi\alpha,\dots,x_{d},y_{1},\dots,y_{d})=(-1)^{f}e^{i\alpha\mu_{E}}K_{S^{1}\times H^{d-1}}(s;x,y),$ (73)
where $\mu_{E}$ represents the purely imaginary chemical potential. The heat kernel $K_{S^{1}}(s;x_{1},y_{1})$ can be expressed as the heat kernel on ${\mathbb{R}}^{1}$:
$K_{\mathbb{R}^{1}}(s;x_{1},y_{1})=\frac{1}{(4\pi s)^{1/2}}e^{\frac{-(x_{1}-y_{1})^{2}}{4s}},$ (74)
using the discrete Sommerfeld formula:
$K_{S^{1}}(s;x_{1},y_{1})=\frac{1}{(4\pi s)^{1/2}}\sum_{m\in\mathbb{Z}}e^{-\frac{(y_{1}-x_{1}+2\pi\alpha m)^{2}}{4s}}e^{(-i\alpha\mu_{E}-i\pi f)m}.$ (75)
Two-Dimensional Case ($d=2$)
For the two-dimensional theory at inverse temperature $\beta_{n}=2\pi n$, the finite-$\mu_{E}$ free energy takes the form:
$F_{n}(\mu_{E})=\frac{(-1)^{f}}{2}(2\pi n)V_{H^{1}}\int\frac{ds}{s}\sum_{m}e^{-\frac{\pi^{2}n^{2}m^{2}}{s}}\frac{1}{\sqrt{4\pi s}}e^{(-in\mu_{E}-i\pi f)m}K_{H^{1}}(\rho=0,s),$ (76)
where $K_{H^{1}}(0,t)=\frac{1}{\sqrt{4\pi t}}$, and $y_{1}-x_{1}=0$ was assumed for the trace. The renormalized free energy eliminates the divergent $m=0$ contribution:
$\hat{F}_{n}(\mu_{E})=\frac{(-1)^{f}}{4}\frac{1}{n\pi^{2}}V_{H^{1}}\sum_{m=1}^{\infty}\frac{e^{i(n\mu_{E}+\pi f)m}+e^{-i(n\mu_{E}+\pi f)m}}{m^{2}}.$ (77)
To compute the free energy using the continuous version of the Sommerfeld formula, we employ the auxiliary function:
$G_{m}(d,\mu_{E})\equiv\frac{e^{-im\mu_{E}}}{m^{d}}.$ (78)
For bosons (i.e., $f=0$), the free energy becomes:
$\hat{F}_{n}(\mu_{E})=\frac{1}{4n\pi^{2}}V_{H^{1}}\sum_{m=1}^{\infty}\left[G_{m}(2,n\mu_{E})+G_{m}(2,-n\mu_{E})\right].$ (79)
Consider the function $\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}$, whose poles are located at $\{m\in\mathbb{Z}\}$. By inserting this function to adjust the position of the poles, we can transform the sum into a contour integral:
$\hat{F}_{n}(\mu_{E})$ $=\frac{1}{4n\pi^{2}}V_{H^{1}}\mathrm{sgn}(\mu_{E})\int_{\Gamma_{2}}\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}G_{m}(2,n\mu_{E}).$ (80)
In this section, the Sommerfeld formula is applied directly to the free energy rather than to the trace of the heat kernel. The two are connected through the integral Eq. (25). The difference in methodology here lies only in the order of integration. Define
$C_{2}(n,\mu_{E})=\mathrm{sgn}(\mu_{E})\int_{\Gamma_{2}}\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}G_{m}(2,n\mu_{E}).$ (81)
This can be expressed by the residue at the origin as:
$C_{2}(n,\mu_{E})$ $=-2\pi i\,\mathrm{sgn}(\mu_{E})\,\mathrm{Res}\left(\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}G_{m}(2,n\mu_{E}),m=0\right)$ (82)
$=\frac{1}{2}\mu_{E}^{2}n^{2}-\pi n|\mu_{E}|+\frac{1}{3}\pi^{2}.$ (83)
Consequently, the free energy and Rényi entropy become:
$\hat{F}_{n}(\mu_{E})$ $=\frac{1}{4n\pi^{2}}V_{H^{1}}\left(\frac{1}{2}\mu_{E}^{2}n^{2}-\pi n|\mu_{E}|+\frac{1}{3}\pi^{2}\right),$ (84)
$S_{n}(\mu_{E})$ $=\left[\frac{1}{12}\left(1+\frac{1}{n}\right)-\frac{|\mu_{E}|}{4\pi}\right]V_{H^{1}}.$ (85)
Four-Dimensional Case ($d=4$)
The four-dimensional extension follows analogous procedures. The modified free energy expression reads:
$\hat{F}_{n}(\mu_{E})$ $=\frac{(-1)^{f}}{2}(2\pi n)V_{H^{3}}\int\frac{ds}{s}\sum_{m\in\mathbb{Z}^{+}}e^{-\frac{\pi^{2}n^{2}m^{2}}{s}}\frac{2\cos((n\mu_{E}+\pi f)m)}{\sqrt{4\pi s}}K_{H^{1}}(0,s)$ (86)
$=\frac{(-1)^{f}}{16}\frac{1}{n^{3}\pi^{5}}V_{H^{3}}\sum_{m=1}^{\infty}\frac{e^{i(n\mu_{E}+\pi f)m}+e^{-i(n\mu_{E}+\pi f)m}}{m^{4}}.$ (87)
Applying the Sommerfeld transformation with $G_{m}(4,\mu_{E})\equiv e^{-im\mu_{E}}/m^{4}$, the free energy and coefficient function become:
$\hat{F}_{n}(\mu_{E})=\frac{1}{16n^{3}\pi^{5}}V_{H^{3}}\mathrm{sgn}(\mu_{E})\int_{\Gamma}\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}G_{m}(4,n\mu_{E}),$ (88)
$C_{2}(n,\mu_{E})$ $=\mathrm{sgn}(\mu_{E})\int_{\Gamma_{2}}\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}G_{m}(4,n\mu_{E})$ (89)
$=-\frac{1}{24}n^{4}\mu_{E}^{4}+\frac{1}{6}\pi n^{3}|\mu_{E}|^{3}-\frac{1}{6}\pi^{2}n^{2}\mu_{E}^{2}+\frac{1}{45}\pi^{4}.$ (90)
This yields the final expressions for free energy and entropy:
$\hat{F}_{n}(\mu_{E})=\left(\frac{|\mu_{E}|^{3}}{96\pi^{4}}+\frac{1}{720\pi n^{3}}-\frac{\mu_{E}^{4}n}{384\pi^{5}}-\frac{\mu_{E}^{2}}{96\pi^{3}n}\right)V_{H^{3}},$ (91)
$S_{n}(\mu_{E})=\left(\frac{(n+1)\mu_{E}^{2}}{48\pi^{2}n}-\frac{|\mu_{E}|^{3}}{48\pi^{3}}-\frac{1+n+n^{2}+n^{3}}{360n^{3}}\right)\frac{V_{H^{3}}}{2\pi}.$ (92)
These calculations demonstrate the systematic application of our improved Sommerfeld formalism to curved backgrounds, reproducing the expected dimensional scaling while incorporating the effects of finite chemical potential. The results provide explicit validation of our framework in holographically relevant geometries.
5 cMERA and Heat Kernel Approach to EE
5.1 From MERA to cMERA
Having established the heat kernel approach for SREE, we now explore its deep connections to cMERA—a tensor network framework that naturally encodes scale-dependent entanglement structures.
The MERA provides a tensor network representation of quantum states that incorporates entanglement renormalization Vidal:2007hda . Its architecture realizes a real-space renormalization group (RG) flow through alternating layers of disentanglers ($U$) and isometries ($W$), which respectively remove short-range correlations and perform coarse-graining. Through this layered structure, the MERA efficiently captures the logarithmic scaling of entanglement in critical systems.
For completeness, the discrete construction using $U$ and $W$ is summarized in Appendix A. In the present section we focus on its continuum analog, where the generators $K(u)$ and $L$ play corresponding roles in quantum field theory.
[…]
$U(u_{2},u_{1})=\mathcal{P}\exp\left({-i}\int_{u_{1}}^{u_{2}}D(u)du\right),\quad D(u)=K(u)+L,$ (93)
where $L$ generates scale transformations, $K(u)$ introduces entanglement across momentum scales, and $\mathcal{P}$ denotes $u$-ordering.
Let $a(k)$ and $a^{\dagger}(k)$ denote annihilation and creation operators in $d$-dimensional momentum space, satisfying:
$[a(k),a^{\dagger}(k^{\prime})]=\delta^{d}(k-k^{\prime}),$ (94)
Under a scale transformation $k\to e^{u}k$, these operators transform as:
$a(k)\to e^{\frac{d}{2}u}a(e^{u}k),\quad a^{\dagger}(k)\to e^{\frac{d}{2}u}a^{\dagger}(e^{u}k).$ (95)
For an infinitesimal transformation $e^{u}=1+\epsilon$, the variation $\delta a(k)=\epsilon(k\partial_{k}+d/2)a(k)$ defines the generator $L$ via:
$-i[L,a(k)]=\left(k\frac{\partial}{\partial k}+\frac{d}{2}\right)a(k).$ (96)
Integrating Eq. (96) yields:
$L=\int d^{d}k\,\left[a^{\dagger}(k)\left(k\frac{\partial}{\partial k}+\frac{d}{2}\right)a(k)+\mathrm{h.c.}\right].$ (97)
A momentum-dependent Bogoliubov rotation mixes $a(k)$ and $a^{\dagger}(-k)$:
$a(k)\to\cosh f(k,u)a(k)+\sinh f(k,u)a^{\dagger}(-k).$ (98)
For infinitesimal $\delta u$, $\delta a(k)=g(k,u)\delta u\,a^{\dagger}(-k)$ , the generator $K(u)$ must satisfy:
$-i[K(u),a(k)]=g(k,u)a^{\dagger}(-k).$ (99)
Then we identify:
$K(u)=\frac{i}{2}\int d^{d}k\,g(u,k)\,[a^{\dagger}(k)a^{\dagger}(-k)-a(-k)a(k)],$ (100)
where $K(u)$ introduces correlations between momentum modes, and $g(u,k)$ parametrizes entanglement strength.
Expressing $L$ and $K(u)$ in terms of $\phi(k)=\frac{1}{\sqrt{2\omega_{k}}}(a(k)+a^{\dagger}(-k))$ and $\pi(k)=-i\sqrt{\frac{\omega_{k}}{2}}(a(k)-a^{\dagger}(-k))$:
$L=\frac{1}{2}\int dk[\pi(-k)(k\partial_{k}+\frac{1}{2})\phi(k)+h.c.],$ (101)
$K(u)=\frac{1}{2}\int dkg(k,u)[\pi(-k)\phi(k)+h.c.].$ (102)
Acting on the field operator $\phi(x)$, we derive:
$-i[D(u),\phi(k)]$ $=-\left(k\partial_{k}+\frac{d}{2}+g(k,u)\right)\phi(k),$ (103)
$-i[D(u),\pi(k)]$ $=-\left(k\partial_{k}+\frac{d}{2}-g(k,u)\right)\pi(k),$ (104)
$U^{-1}(0,u)\phi(k)U(0,u)$ $=e^{-f(k,u)}e^{-\frac{d}{2}u}\phi(e^{-u}k),$ (105)
$U^{-1}(0,u)\pi(k)U(0,u)$ $=e^{f(k,u)}e^{-\frac{d}{2}u}\pi(e^{-u}k),$ (106)
where
$f(k,u)=\int_{0}^{u}g(ke^{-s},s)ds.$ (107)
The entanglement entropy is generally determined by both the geometry of the entangling surface $\Sigma$ and its embedding in ambient spacetime. However, in certain simple cases, the entropy depends solely on the intrinsic geometry of $\Sigma$.
5.2 SREE in Gaussian cMERA
As established in Section 3.2.2 for neutral systems, the half-space entanglement entropy depends solely on the intrinsic geometry of the entangling surface $\Sigma$. This geometric insight can be naturally extended to Gaussian cMERA states.
Using the effective action $W(\alpha)$, the entropy can be written as
$S=\left(\alpha\partial_{\alpha}-1\right)W(\alpha)\big|_{\alpha=1},\quad W(\alpha)=\sum_{i=0}^{\infty}w_{i}(1-\alpha)^{i},$ (108)
so that
$S=-(w_{0}+w_{1}).$ (109)
For a $d$-dimensional planar surface of area $V$ in a neutral theory, the heat kernel trace exhibits the following universal form:
$\operatorname{Tr}K(s)=\frac{V}{(4\pi s)^{d/2}}.$ (110)
As established in Eq. (55), when extended to the replicated manifold via the Sommerfeld formula, this transforms to:
$\operatorname{Tr}K_{\alpha}(s)$ $=\alpha\operatorname{Tr}K(s)+\operatorname{Tr}\Delta_{\alpha}$ (111)
$=\frac{1}{(4\pi s)^{d/2}}V\cdot\alpha+\frac{1}{(4\pi s)^{(d-2)/2}}A(\Sigma)\cdot\frac{\alpha C_{2}(\alpha)}{2}$ (112)
$=\alpha\operatorname{Tr}K(s)+\frac{\alpha C_{2}(\alpha)}{2}\operatorname{Tr}K_{\Sigma}(s),$ (113)
where the surface term $\operatorname{Tr}\Delta_{\alpha}$ originates from the (d2)(d-2)-dimensional entangling surface $\Sigma$ Callan:1994py , $K_{\Sigma}$ denotes the trace of the heat kernel on $\Sigma$, and the geometric coefficient $C_{2}(\alpha)$ is derived in Eq. (53):
$\alpha C_{2}(\alpha)=\frac{1}{6}\left(\frac{1}{\alpha}-\alpha\right)\sim\frac{1}{3}(1-\alpha)+O\left((\alpha-1)^{2}\right).$ (114)
Expanding $\operatorname{Tr}\Delta_{\alpha}$ near $\alpha=1$ gives the leading contribution relevant for entanglement:
$\operatorname{Tr}\Delta_{\alpha}=\operatorname{Tr}K_{\Sigma}(s)\cdot\frac{1}{6}\cdot(1-\alpha).$ (115)
Integrating over the proper time $s$,
$w_{0}=0,\quad w_{1}=-\frac{1}{2}\int\frac{ds}{s}\operatorname{Tr}K_{\Sigma}(s)\cdot\frac{1}{6},$ (116)
The formalism naturally extends to charged quantum fields through the modified Sommerfeld prescription developed in Section 3. For a system with chemical potential $\mu=i\mu_{E}$, the Sommerfeld-corrected heat kernel trace reads
$\operatorname{Tr}K_{\alpha\mu}(s)=\alpha\operatorname{Tr}K(s)+\operatorname{Tr}\Delta_{\alpha\mu}=\alpha\operatorname{Tr}K(s)+\frac{\alpha C_{2}(\alpha,\mu_{E})}{2}\operatorname{Tr}K_{\Sigma}(s),$ (117)
where the coefficient function $C_{2}(\alpha,\mu_{E})$ from Eq. (64) takes the form
$C_{2}(\alpha,\mu_{E})=\frac{1}{6}\left(\frac{1}{\alpha^{2}}-1\right)+\frac{\mu_{E}^{2}}{4\pi^{2}\alpha^{2}}-\frac{|\mu_{E}|}{2\pi\alpha^{2}}.$ (118)
Expanding near $\alpha=1$ gives the correction in the effective action:
$\alpha\left[\frac{1}{6}\left(\frac{1}{\alpha^{2}}-1\right)+\frac{\mu_{E}^{2}}{4\pi^{2}}-\frac{|\mu_{E}|}{2\pi\alpha}\right]\sim\frac{\mu_{E}^{2}-2\pi|\mu_{E}|}{4\pi^{2}}-\left(\frac{|\mu_{E}|}{2\pi}-\frac{\mu_{E}^{2}}{4\pi^{2}}-\frac{1}{3}\right)(1-\alpha)+O\left((1-\alpha)^{2}\right).$ (119)
This yields
$w_{1}+w_{0}=-\frac{1}{2}\int\frac{ds}{s}\operatorname{Tr}K_{\Sigma}(s)\cdot\left(\frac{1}{6}-\frac{|\mu_{E}|}{2\pi}+\frac{\mu_{E}^{2}}{4\pi^{2}}\right).$ (120)
the entanglement entropy is expressed as
$S=\frac{1}{12}\int_{\epsilon^{2}}^{\infty}\frac{ds}{s}\operatorname{Tr}K_{\Sigma}(s)=-\frac{1}{12}\ln\det(-\Delta(\Sigma)),$ (121)
and the charged entanglement entropy becomes
$S(\mu_{E})$ $=\left(\frac{1}{12}-\frac{|\mu_{E}|}{4\pi}+\frac{\mu_{E}^{2}}{8\pi^{2}}\right)\int_{\epsilon^{2}}^{\infty}\frac{ds}{s}\operatorname{Tr}K_{\Sigma}(s)$ (122)
$=-\left(\frac{1}{12}-\frac{|\mu_{E}|}{4\pi}+\frac{\mu_{E}^{2}}{8\pi^{2}}\right)\ln\det(-\Delta(\Sigma)).$ (123)
where $\Delta(\Sigma)$ is the Laplacian operator defined on the (d2)(d-2)-dimensional surface $\Sigma$. The entropy can then be expressed in terms of the Green function as
[…]
$a_{k}=\sqrt{\frac{\alpha(k)}{2}}\phi(k)+i\sqrt{\frac{1}{2\alpha(k)}}\pi(k),$ (125)
[…]
The commutation relation $\phi(k),\pi(p)]=i\delta(k+p)$ governs the evaluation of expectation values. Substituting the above relations into this structure, one finds:
$\braket{\pi(k)\pi(p)}$ $=\frac{\alpha(k)}{2}\delta(k+p),$ $\braket{\phi(k)\phi(p)}$ $=\frac{1}{2\alpha(k)}\delta(k+p),$ $\braket{\phi(k)\pi(p)}$ $=\frac{i}{2}\delta(k+p).$ (127)
[…]
$U^{-1}(0,-\infty)\phi(k)U(0,-\infty)=e^{-f(k,-\infty)}\phi(k),$ (128)
the two-point function in the UV state reads
[…]
Thus, the entropy can be expressed in terms of the functions in cMERA as
$S=-\frac{A(\Sigma)}{3}\int d^{d-1}kf(k,\infty)+C^{\prime}.$ (130)
Introducing $g(k,u)=g(u)\Gamma(k/\Lambda)$ and defining the cMERA weight function
$\Sigma(u)=\int d^{d-1}k\,k^{d-2}\Gamma(ke^{-u}/\Lambda),$ (131)
we arrive at the differential form Fernandez-Melgarejo:2021ymz
$\frac{dS}{du}=-\frac{A(\Sigma)}{3}\cdot g(u)\cdot\Sigma(u).$ (132)
Correspondingly, the cMERA charged entanglement flow equation becomes
$\frac{dS(\mu_{E})}{du}=-\left(\frac{1}{3}-\frac{|\mu_{E}|}{\pi}+\frac{\mu_{E}^{2}}{2\pi^{2}}\right)A(\Sigma)g(u)\Sigma(u),$ (133)
which generalizes the neutral case and explicitly shows the suppression of entropy growth due to the chemical potential. This result demonstrates how the background charge deforms the entanglement structure at each renormalization group scale, which is directly reflected in the modified profile of the entangler function $g(u;\mu_{E})$. The expression provides a direct bridge between charged field theory and geometric entropy within the holographically motivated Gaussian cMERA framework, explicitly showing how the chemical potential suppresses entropy growth at all scales.
6 Conclusions
The improved Sommerfeld formula developed in this work resolves challenges in computing entanglement entropy for charged quantum fields. By incorporating a phase factor $\frac{1}{1-e^{-2\pi iw/\alpha\,\mathrm{sgn}[\mu]}}$ into the kernel function, we systematically handle the singularities induced by chemical potentials while preserving the geometric intuition of the replica trick. This modification ensures rigorous handling of contour integrals and residue contributions, enabling precise calculations of symmetry-resolved entropies in both flat and curved spacetimes.
Our key contributions include three main results. First, the derived entropy formulas generalize across spacetime dimensions, reducing to the known logarithmic scaling in $d=2$ Calabrese:2004eu ; Calabrese:2009qy ; Calabrese:2016xau and agreeing with holographic predictions for $S^{1}\times H^{d-1}$ backgrounds Belin:2013uta . Second, independent validations against twist operator correlators in $(1+1)D$ CFT Murciano:2020vgh and free energy calculations in higher-dimensional AdS/CFT setups Belin:2013uta ; Huang:2025lsy confirm the robustness of our method. Third, extending the known cMERA representation of neutral entanglement entropy Fernandez-Melgarejo:2021ymz , we establish a heat-kernel/cMERA correspondence for charged systems. We derive a modified entanglement flow equation in the presence of a finite chemical potential and show that the Gaussian cMERA formalism naturally captures the microscopic mechanism of symmetry resolution across renormalization scales.
Thus, the improved heat kernel method provides a powerful geometric tool for probing symmetry-resolved entanglement, unifying the treatment of charged and neutral sectors, and offering a versatile approach applicable to both conformal field theories and their holographic duals.
Although the heat-kernel method successfully establishes a link between cMERA and SREE – achieved here through calculation correlation functions within the cMERA framework and their subsequent use in the heat-kernel expression for entropy – this connection remains indirect. It relies on the heat kernel as an intermediary computational tool that interprets cMERA-generated data into entanglement measures Fernandez-Melgarejo:2021ymz .
A profound challenge and a central direction for future work is therefore to develop a direct cMERA formulation of SREE. This would move beyond the current dependence on auxiliary field-theoretic machinery, constructing instead a symmetry-adapted cMERA where $U(1)$ charge conservation is manifestly encoded at every scale $u$ in the renormalization group flow. Key objectives include developing techniques to extract charged moments $Z_{n}(q)$ or the resolved density matrix with symmetry $\rho_{A}(q)$ directly from the cMERA state, potentially by leveraging the interaction between the disentangler $K(u)$ and the emergent geometry. This would bypass the need for both the heat kernel and replica trick in the SREE calculations for these states.
Such a direct cMERA-SREE framework could be utilized to study how symmetry resolution evolves under the renormalization group. Key questions include: Does equipartition Xavier:2018kqb emerge universally along the flow towards critical points? Moreover, how does the fine-grained structure of symmetry sectors in the boundary cMERA state relate to specific geometric features (e.g., charged minimal surfaces, flux threads, or defect structures) in the bulk gravitational dual? A direct cMERA approach to SREE promises a more intrinsic understanding of the holographic dictionary for symmetry-resolved entanglement. Establishing such connections would provide a richer microscopic picture of how gauge symmetries and entanglement intertwine in quantum gravity Swingle:2009bg .
The pursuit of this direct connection represents more than a technical refinement – it seeks to uncover the intrinsic entanglement structures woven into the renormalization group flow itself. Success would solidify cMERA not just as an efficient variational tool but as a fundamental framework capable of directly revealing how global symmetries organize and constrain quantum entanglement across scales – a question of deep significance for quantum field theory, quantum information, and our understanding of holography.
Our calculation of SREE in cMERA can guide the computation in discrete MERA tensor network state. Our result Eq. (133) imply that the charge reduces the bond dimension of the charge sector of the tensors along the minimal surface in the MERA.
Finally, our calculation of SREE within the cMERA formalism provides valuable insight for the discrete MERA tensor network. Equation (133) imply that $U(1)$ charge effectively reduce the bond dimension of the charge sector tensors along the minimal surface in MERA.
Acknowledgements.
We thank Giuseppe Di Giulio and Suting Zhao for helpful discussions. ZYX acknowledges funding from the DFG through the Collaborative Research Center SFB 1170 “ToCoTronics” (Project ID 258499086 – SFB 1170) and support from the Berlin Quantum Initiative.
Appendix A Discrete MERA construction
For completeness, we summarize here the discrete formulation of the multi-scale entanglement renormalization ansatz (MERA), which provides the operational intuition behind the continuum version (cMERA) discussed in Sec. 5. The discrete MERA is built from alternating layers of disentanglers $W$ and isometries $U$, which act locally to remove short-range correlations and perform coarse-graining in the Hilbert space. Their structure follows from the basic principles of bipartite entanglement and Schmidt decomposition, as reviewed below.
For a Hilbert space decomposition $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ with $\dim\mathcal{H}_{A}=\dim\mathcal{H}_{B}=d$, any pure state admits the canonical Schmidt form:
$|\psi\rangle=\sum_{\alpha=1}^{m}\lambda_{\alpha}|\alpha\rangle_{A}\otimes|\alpha\rangle_{B},\quad\lambda_{1}\geq\lambda_{2}\geq\cdots>0,\quad\sum_{\alpha=1}^{m}\lambda_{\alpha}^{2}=1,$ (A.1)
where the Schmidt rank $m\leq d$ quantifies bipartite entanglement. A local unitary operator $U:\mathbb{C}^{d}\otimes\mathbb{C}^{d}\to\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ can be chosen to rotate local bases and concentrate entanglement into a smaller number of Schmidt modes:
$U|\psi\rangle=\sum_{\beta=1}^{\tilde{m}}\tilde{\lambda}_{\beta}|\tilde{\beta}\rangle_{A}\otimes|\tilde{\beta}\rangle_{B},\quad\tilde{m}\ll m.$ (A.2)
This step corresponds to the action of a disentangler in the MERA layer.
The subsequent isometry $W:\mathbb{C}^{d}\otimes\mathbb{C}^{d}\to\mathbb{C}^{\chi}$ ($\chi\leq d^{2}$) maps the two-site Hilbert space to a reduced effective space:
$W=\sum_{k=1}^{\chi}|k\rangle_{C}\,(\langle\tilde{k}|_{A}\otimes\langle\tilde{k}|_{B}),$ (A.3)
yielding a compressed state,
$WU|\psi\rangle=\sum_{k=1}^{\chi}\tilde{\lambda}_{k}|k\rangle_{C},$ (A.4)
with truncation error
$\epsilon=\sum_{k=\chi+1}^{d^{2}}\tilde{\lambda}_{k}^{2}.$ (A.5)
The full MERA tensor network arises from iterative application of the disentangler–isometry pair across multiple layers:
$|\psi^{(\ell+1)}\rangle=W^{(\ell)}U^{(\ell)}|\psi^{(\ell)}\rangle,\quad\ell=0,\ldots,L-1,$ (A.6)
where $\mathcal{H}^{(\ell)}$ denotes the Hilbert space at depth $\ell$. The sequence of bond dimensions $\{\chi_{\ell}\}$ decreases monotonically,
$\chi_{0}\geq\chi_{1}\geq\cdots\geq\chi_{L}\geq 1,$ (A.7)
encoding the gradual loss of microscopic degrees of freedom along the renormalization direction.